Suppose that the economy is characterized by the following equations: C = 160 + 0.6Yd I = 150; G = 150; and T = 100: For the economy in the last problem above:

(a) Assume that output is equal to 900. Compute total demand. Is it equal to output? Explain.

(b) Assume that output is equal to 1000. Compute total demand. Is it equal to output? Explain.

(c) Assume output is equal to 1000. Compute private saving. Is it equal to investment? Explain
Answer #1

Total demand = C + I + G

Yd = Y - T

a) Total demand = 160 + 0.6 x (900 - 100) + 150 + 150 = 940

No demand is not equal to output, Demand exceeds the output (there would be shortfall of inventory)

b) Total demand = 160 + 0.6 x (1000 - 100) + 150 + 150 = 1000

Demand is equal to output, Demand equals the output (equilibrium)

c) No private saving is not equal to investment

Yd = C + S

S = Yd - C = 1000 - 100 - 160 - 0.6 x (1000 - 100) = 200